# Multi-Channel Hypergraph Contrastive Learning for Matrix Completion

### v1

@@ 5.1.1. Datasets
To assess the efficacy of MHCL, we conduct matrix completion and recommendation experiments on 6 publicly available datasets, including Yelp (Seo et al., 2017), Amazon (Ni et al., 2019), MovieLens-100K (ML-100K for short) (Miller et al., 2003), MovieLens-1M (ML-1M for short) (Miller et al., 2003), YahooMusic (Dror et al., 2012) and Douban (Miller et al., 2003). The summary of dataset statistics is presented in Table 1. A detailed description of datasets can be found in Sec. A.2 of the Supplement.
@@
Yelp 8423 3742 88647 0.281 5.88 9.41 15.29 37.13 32.09
Amazon 5541 3568 64706 0.330 4.31 4.65 10.49 25.55 54.98
ML-100K 943 1682 100000 6.30 6.11 11.37 27.14 34.17 21.2
ML-1M 6040 3706 1000209 4.47 5.61 10.75 26.12 34.89 22.63
YahooMusic 3000 3000 5335 0.06 1.04 11.84 19.15 19.03 48.94
Douban 3000 3000 136891 1.52 5.87 11.25 25.71 30.23 26.94
@@ 5.1.2. Baselines
To demonstrate the effectiveness, we compare our proposed MHCL with the following four categories of methods: (1) traditional methods: GRALS (Rao et al., 2015), NNMF (Dziugaite and Roy, 2015); (2) GAE-based methods: GC-MC (Berg et al., 2017), IMC-GAE (Shen et al., 2021); (3) inductive methods: PinSage (Ying et al., 2018), IGMC (Zhang and Chen, 2019), GIMC (Zhang et al., 2022), IDCF-NN (Wu et al., 2021b) and IDCF-GC (Wu et al., 2021b); (4) review-based methods: SSG (Gao et al., 2020a), RGCL (Shuai et al., 2022) and MAGCL (Wang et al., 2023). Besides, we also introduce: (5) review and rating-based recommendation methods: SGL (Wu et al., 2021a), DGCLR (Ren et al., 2022), VGCL (Yang et al., 2023), and ReHCL (Wang et al., 2024) for comparison to evaluate the recommendation performance of MHCL. A detailed description of baselines can be found in Sec. A.3 of the Supplement.
@@ 5.1.3. Evaluation Metrics and Experimental Settings
We use MSE and MAE metrics for evaluating the matrix completion task, and MRR@10 (M@10 for short) and NDCG@10 (N@10 for short) for the recommendation task. The experimental settings are detailed in the Sec. A.4 of the Supplement.
@@ 5.2. Matrix Completion (RQ1)
As shown in Table 2, MHCL consistently outperforms all baselines across datasets, with the most +9.04% notable MSE improvement on ML-100K. MHCL excels without needing extra side information or review data, making it simpler than models requiring complex preprocessing. It also surpasses review-based models on Amazon and Yelp datasets, and is more efficient than subgraph-based models, avoiding the significant time and memory costs associated with creating individual subgraphs for each rating edge, especially in large datasets. The performance highlights MHCL’s superiority in mitigating data sparsity and long-tail distribution issues.
@@
Traditional matrix completion methods such as GRALS and NNMF perform well on small datasets but struggle with scalability, while GNN-based models excel in this area. Although GC-MC and IMC-GAE provide fast runtimes, they converge more slowly and deliver slightly inferior performance compared to MHCL. Inductive models like IGMC and GIMC achieve competitive results but require substantial time investment, with their performance on certain datasets not matching that of MHCL. Models such as PinSage, IDCF-NN, and IDCF-GC effectively capture local patterns but fail to model higher-order dependencies and cross-rating features, which diminishes their accuracy in ordinal matrix completion tasks. Review-based models like SSG and RGCL leverage user comments to enhance embeddings but cannot function without review data. In contrast, MHCL dynamically learns hypergraph structures for each rating subgraph, capturing both local and global dependencies, and integrates these representations through cross-rating fusion. Notably, MHCL does not rely on review data yet consistently outperforms review-based methods on datasets like Amazon and Yelp, showcasing its versatility. Moreover, MHCL offers significant advantages in time complexity, making it a scalable and efficient solution for large-scale matrix completion tasks.
@@ 5.3. Recommendation (RQ2)
Next, we evaluate the recommendation performance by comparing MHCL with the latest recommendation baselines which consider user rating matrices. Expect baselines in Table 2, we add the review and rating-based recommendation baselines (SGL, DGCLR, VGCL, ReHCL). Results are reported in Table 3, MHCL consistently outperforms all advanced baselines across datasets, with the most +10.31% notable N@10 improvement on ML-100K and average +6.65% improvement on M@10 and N@10 across all datasets. The performance highlights MHCL’s superiority in recommendation tasks.
@@
In recommendation tasks, many matrix-completion methods integrate user rating matrices into node representations to effectively capture user preferences and enhance overall recommendation performance. For instance, on Amazon and Yelp, which include user reviews, the current state-of-the-art method, ReHCL, leverages both the rating matrix and the textual reviews provided by users. In contrast, our method, MHCL, achieves an average performance improvement of +4.14% over ReHCL without incorporating any review data, relying solely on the rating matrix. This outcome highlights the effectiveness and necessity of our matrix completion algorithm design, demonstrating that a well-structured rating matrix can be sufficient for improving recommendation quality. Furthermore, when analyzing four datasets that do not contain reviews, we observe a maximum performance improvement of over 10% in both M@10 and N@10 metrics compared to IDCF-GC, the leading matrix completion method among the baselines. This finding suggests that the rating matrices generated through MHCL not only capture more nuanced user preference information but also maintain their effectiveness across diverse datasets. Overall, MHCL’s ability to deliver superior performance without relying on review data underscores its potential as a robust solution for matrix completion in recommendation systems.
5.4. Ablation Study (RQ3)
@@
To evaluate the effectiveness of each component of our model, we have further conducted an ablation study on different MHCL variations. We report the results of the ablation study on five datasets in Table 4. Specifically, we generate the following variants:
@@
  • w/o NRR - which removes the $\mathcal{L}_{NRR}$ in the final loss function.
@@
Our results yield several key insights. First, removing the cross-ratings representation fusion (CRF) module (w/o CRF) leads to performance degradation across all datasets, underscoring its critical importance in capturing complex user interactions. Similarly, omitting the multi-channel cross-ratings module and global hypergraph learning (w/o C-R and w/o Glo) also results in reduced performance. Notably, the cross-ratings module significantly enhances results for Amazon, where the rating distribution is imbalanced, while hypergraph learning offers particular benefits for denser datasets such as ML-100K and ML-1M. The variant without contrastive learning (w/o CL) demonstrates the poorest performance, highlighting the crucial role of contrastive learning in enriching node representations and improving model robustness. Additionally, removing the balanced cross-entropy loss (w/o bCE) adversely affects performance, further confirming its effectiveness in capturing correlations between closely-rated scores. Furthermore, our analysis indicates that the cross-ratings module (C-R) has the most substantial impact on the loss function, whereas node relation regularization (NRR) exhibits the least effect, aligning with our parameter sensitivity findings. Finally, while the ordinal cross-entropy loss (with OCE) performs similarly to the balanced cross-entropy loss, the latter is preferred due to its slight performance edge and superior computational efficiency, making it a more practical choice for implementation.
Table 4. Performance comparison of different variants (MSE).
Dataset Amazon YELP ML-100K ML-1M YahooMusic
MHCL (w/o C-R) 0.7839 1.1320 0.7856 0.6842 336.72
MHCL (w/o Glo) 0.7703 1.1354 0.7899 0.6874 340.53
MHCL (w/o CL) 0.7884 1.1392 0.7910 0.6905 350.77
MHCL (w/o CRF) 0.7701 1.1320 0.7857 0.6857 342.13
MHCL (w/o bCE) 0.7722 1.1343 0.7836 0.6851 339.85
MHCL (w/o NRR) 0.7689 1.1328 0.7845 0.6869 340.05
MHCL (with OCE) 0.7651 1.1298 0.7796 0.6799 335.95
MHCL 0.7610 1.1289 0.7790 0.6790 334.89
5.5. Impact of imbalanced data distribution (RQ4)
@@ 5.5.1. Impact of Imbalanced User Interactions
We evaluate MHCL’s performance across users with varying interaction sparsity by dividing them into three groups: Inactive (bottom 80% by interaction frequency), Normal (middle 15%), and Active (top 5%). We then compare MHCL with two leading baselines on two datasets, with results shown in Figures LABEL:fig2.1 and LABEL:fig2.2. MHCL consistently outperforms the baselines, with the largest gains observed among Inactive users. This improvement is driven by InfoNCE loss, which mitigates popularity bias by enhancing representation uniformity.
@@ 5.5.2. Impact of Imbalanced Rating Interactions
Experimental results according to different rating categories are shown in Figures LABEL:fig2.3 and LABEL:fig2.4. The findings highlight significant improvements in long-tail ratings (‘1’ and ‘2’) without sacrificing performance in the head classes. This underscores the effectiveness of InfoNCE in reshaping embedding distributions to better capture long-tail characteristics. The impact is especially notable in imbalanced datasets like Amazon, where ratings ‘1’ and ‘2’ constitute less than 10%, making MHCL’s enhancements particularly striking.
5.6. Parameter Sensitivity (RQ5)
5.6.1. Effect of Hyperparameter $\alpha$, $\beta$ and $\lambda$.
We analyze the sensitivity of parameters $\alpha$, $\beta$, and $\lambda$ on contrastive learning (CL) losses using the Amazon dataset (Fig. LABEL:alpha_amazon, LABEL:beta_amazon, LABEL:lambda_amazon). Increasing these parameters from 0 to 0.01 enhances MHCL’s performance, but larger values can impair the main loss function. The optimal range of 0.01 to 0.05 highlights the need to balance CL with the main loss for optimal results.

### v2

@@ 5.1.1. Datasets
To assess the efficacy of MHCL, we conduct matrix completion and recommendation experiments on eight publicly available datasets, including Yelp (Seo et al., 2017), Amazon (Ni et al., 2019), MovieLens-100K (ML-100K for short) (Miller et al., 2003), MovieLens-1M (ML-1M for short) (Miller et al., 2003), YahooMusic (Dror et al., 2012), Douban (Miller et al., 2003), CDs and Vinyl (CD for short) (Wang et al., 2023), and Alibaba-iFashion (Alibaba for short) (Chen et al., 2019). The summary of dataset statistics is presented in Table 2. For YahooMusic dataset, we utilize preprocessed subsets of this dataset as provided by (Monti et al., 2017). Amazon and Yelp datasets are distinct in that they include not only user-item ratings but also detailed review information. These datasets encompass four primary data components: users, items, ratings, and review documents. We focus on the ‘Digital Music’ domain within the Amazon dataset, specifically the Amazon-5core dataset in our experiments, where ‘5-core’ indicates that each user or item has a minimum of five reviews. Yelp dataset originates from the Yelp Business Rating Prediction Challenge 2013 dataset, which comprises restaurant reviews from the Phoenix, AZ metropolitan area. Following the preprocessing steps outlined in (Shuai et al., 2022), we establish a 5-core dataset from the raw Yelp data. Alibaba is a sparse dataset, and we randomly sample 300k users and use all their interactions with fashion outfits.
@@
Yelp 8,423 3,742 88,647 0.281 5.88 9.41 15.29 37.13 32.09
Amazon 5,541 3,568 64,706 0.330 4.31 4.65 10.49 25.55 54.98
ML-100K 943 1,682 100,000 6.30 6.11 11.37 27.14 34.17 21.2
ML-1M 6,040 3,706 1,000,209 4.47 5.61 10.75 26.12 34.89 22.63
YahooMusic 3,000 3,000 5,335 0.06 1.04 11.84 19.15 19.03 48.94
Douban 3,000 3,000 136,891 1.52 5.87 11.25 25.71 30.23 26.94
CD 75,258 64,443 1,097,592 0.023 5.11 11.25 23.26 29.77 30.61
Alibaba 300,000 81,614 1,607,813 0.007 4.43 9.84 27.08 30.50 28.15
@@ 5.1.2. Baselines
To demonstrate the effectiveness, we compare our proposed MHCL with the following four categories of methods: (1) traditional methods: GRALS (Rao et al., 2015), NNMF (Dziugaite and Roy, 2015); (2) GAE-based methods: GC-MC (Berg et al., 2017), IMC-GAE (Shen et al., 2021); (3) inductive methods: PinSage (Ying et al., 2018), IGMC (Zhang and Chen, 2019), GIMC (Zhang et al., 2022), IDCF-NN (Wu et al., 2021b) and IDCF-GC (Wu et al., 2021b); (4) rating-based methods: SGL (Wu et al., 2021a), and VGCL (Yang et al., 2023). Besides, we also introduce: (5) rating and review-based recommendation methods: SSG (Gao et al., 2020a), RGCL (Shuai et al., 2022),MAGCL (Wang et al., 2023),DGCLR (Ren et al., 2022), and ReHCL (Wang et al., 2024) for comparison to evaluate the recommendation performance of MHCL.
Traditional Method:
  • NNMF (Dziugaite and Roy, 2015) is a classical matrix factorization model which uses the user-item ratings (or interactions) only as the target value of its objective function. It explores the possibility of substituting the inner product with a learned function, concurrently with the learning process of latent feature vectors.
GAE-based Methods:
  • GC-MC (Berg et al., 2017) regards rating prediction as link prediction on the user-item bipartite graph and adopts Graph Autoencoder (GAE) to encode user and item embeddings and proposes a graph auto-encoder framework that can utilize differentiable message passing on the bipartite interaction graph.
Inductive Methods:
  • PinSage (Ying et al., 2018) combines a highly effective random walk and graph convolution to represent node embeddings that incorporate both node features and graph structure. It also introduces an innovative training strategy, which progressively incorporates more challenging training examples to enhance the model’s robustness and convergence.
Rating-based Methods:
Rating and Review-based Methods:
  • SSG (Gao et al., 2020a) introduces a multi-view approach to enhance existing single-view methods, incorporating two additional perspectives: sequence and graph, for leveraging reviews. Specifically, the three-way encoder architecture captures long-term (set), short-term (sequence), and collaborative (graph) user-item features for recommendations.
It is worth noting that certain methods within these groups require specific preprocessing steps before model training. For instance, RGCL employs BERT-Whitening (Su et al., 2021) to convert textual comments into vector representations, while both IGMC and GIMC mandate the extraction of subgraphs for all edges within the dataset to operate effectively. For most of the baselines, we split our training, validation, and test sets in the ratio of 8:1:1, following the same settings in recent studies (Zhang et al., 2022; Shen et al., 2021).
@@ 5.1.3. Evaluation Metrics and Experimental Settings
We use MSE and MAE metrics for evaluating the matrix completion task, and MRR@10 (M@10 for short) and NDCG@10 (N@10 for short) for the recommendation task.
[…]
Specifically, in the recommendation section, we utilize the following loss function to evaluate the model performance:
@@ 5.2. Matrix Completion (RQ1)
As shown in Table 3, MHCL consistently outperforms all baselines across diverse datasets, with the most notable +9.04% MSE improvement on ML-100K. MHCL excels without needing extra side information or review data, making it simpler than models requiring complex preprocessing. It also surpasses review-based models on Amazon and Yelp datasets, avoiding the significant time and memory costs associated with creating individual subgraphs for each rating edge, especially in large datasets. In the other four datasets without review information (ML-100K, ML-1M, YahooMusic, and Douban), compared to the suboptimal methods IMC-GAE and GIMC, MHCL demonstrates substantial improvements, achieving a maximum MAE performance improvement of +4.34% on YahooMusic dataset. The performance highlights MHCL’s superiority in mitigating data sparsity and long-tail distribution issues.
@@
Traditional matrix completion methods, such as GRALS and NNMF, perform well on small datasets but struggle with scalability, while GNN-based models excel in this area. Although GC-MC and IMC-GAE provide fast runtimes, they converge more slowly and deliver slightly inferior performance compared to MHCL. Inductive models like IGMC and GIMC achieve competitive results but require substantial time investment, with their performance on certain datasets not matching that of MHCL. Models such as PinSage, IDCF-NN, and IDCF-GC effectively capture local patterns but fail to model higher-order dependencies and cross-rating features, which diminishes their accuracy in ordinal matrix completion tasks. Review-based models like SSG and RGCL leverage user comments to enhance embeddings but cannot function without review data. In contrast, MHCL dynamically learns hypergraph structures for each rating subgraph, capturing both local and global dependencies, and integrates these representations through cross-rating fusion. Notably, MHCL does not rely on review data yet consistently outperforms review-based methods on datasets like Amazon and Yelp, showcasing its versatility. Moreover, MHCL offers significant advantages in time complexity, making it a scalable and efficient solution for large-scale matrix completion tasks.
@@ 5.3. Recommendation (RQ2)
Next, we evaluate the recommendation performance by comparing MHCL with the latest recommendation baselines which consider user rating matrices. Expect baselines in Table 3, we add the review and rating-based recommendation baselines (SGL, DGCLR, VGCL, ReHCL). Experimental results are reported in Table 4, MHCL consistently outperforms all advanced baselines across datasets, with the most notable +10.31% N@10 improvement on ML-100K and average +6.65% improvement on M@10 and N@10 across all datasets. The performance highlights MHCL’s superiority in recommendation tasks.
@@
In recommendation tasks, many matrix-completion methods integrate user rating matrices into node representations to effectively capture user preferences and enhance overall recommendation performance. For instance, on Amazon and Yelp, which include user reviews, the current state-of-the-art method, ReHCL, leverages both the rating matrix and the textual reviews provided by users. In contrast, our method, MHCL, achieves an average performance improvement of +4.14% over ReHCL without incorporating any review data, relying solely on the rating matrix. This outcome highlights the effectiveness and necessity of our matrix completion algorithm design, demonstrating that a well-structured rating matrix can be sufficient for improving recommendation quality. Furthermore, when analyzing four datasets that do not contain reviews, we observe an average performance improvement of +10.20% in M@10 and N@10 metrics on ML-100K dataset compared to IDCF-GC, the leading matrix completion method among the baselines. This finding suggests that the rating matrices generated through MHCL not only capture more nuanced user preference information but also maintain their effectiveness across diverse datasets. Overall, MHCL’s ability to deliver superior performance without relying on review data underscores its potential as a robust solution for matrix completion in recommender systems.
5.4. Performance on large-scale datasets (RQ3)
Table 5. Performance comparison of models on large-scale graph datasets.
Method CD Alibaba
Matrix Completion Recommendation Matrix Completion Recommendation
MSE MAE M@10 N@10 MSE MAE M@10 N@10
GRALS 0.8558 0.6980 0.3509 0.4326 2.0459 1.4434 0.1049 0.0460
NNMF 0.8640 0.7077 0.3515 0.4340 2.0547 1.4550 0.1054 0.0466
GC-MC 0.8404 0.6828 0.3533 0.4351 2.0305 1.4327 0.1070 0.0485
IMC-GAE 0.8234 0.6796 0.3546 0.4368 2.0140 1.4156 0.1069 0.0486
PinSage 0.8726 0.7151 0.3555 0.4382 2.0613 1.4620 0.1081 0.0498
IGMC 0.8215 0.6767 0.3559 0.4388 OOM. OOM. OOM. OOM.
GIMC 0.8210 0.6759 0.3579 0.4403 OOM. OOM. OOM. OOM.
IDCF-NN 0.8201 0.6751 0.3597 0.4415 2.0101 1.4119 0.1095 0.0514
IDCF-GC 0.8193 0.6725 0.3607 0.4450 2.0089 1.4109 0.1101 0.0517
SGL 0.8151 0.6691 0.3685 0.4452 2.0079 1.4090 0.1118 0.0525
VGCL 0.8121 0.6654 0.4030 0.4786 2.0066 1.4075 0.1131 0.0540
[…]
MHCL 0.7655 0.6259 0.4471 0.5280 1.9789 1.3858 0.1239 0.0595
To evaluate MHCL on large-scale datasets CD and Alibaba, we compare it with SOTA methods on both matrix completion and recommendation tasks. Notice that review-based methods do not directly apply to the Alibaba dataset, as they do not contain review data. To facilitate a fair comparison, SGL, DGCLR, VGCL, and ReHCL, originally designed for recommendation, have been adapted to the matrix completion task using a unified loss function and evaluation metrics. Results in Table 5 highlight the clear superiority of MHCL over all baselines. Specifically, in the matrix completion task, MHCL achieves up to +5.48% improvement on two datasets, demonstrating its enhanced ability to model complex relationships between users and items while maintaining model robustness. In the recommendation task, MHCL significantly outperforms rating and review-based recommendation methods, marking a substantial average +8.61% and +13.81% improvement on CD and Alibaba datasets, respectively. These results underscore MHCL’s effectiveness in capturing fine-grained dependencies while scaling efficiently, making it capable of being applied to large-scale graphs.
5.5. Ablation Study (RQ4)
@@
To evaluate the effectiveness of each component of our model, we have further conducted an ablation study on different MHCL variations. We report the results of the ablation study on four datasets in Figure LABEL:fig:ablation, with MSE metric in the matrix completion task and NDCG@10 metric in the recommendation task. Specifically, we generate the following variants:
@@
  • w/o NRR - which removes the $\lambda\mathcal{L}_{NRR}$ in the final loss function.
@@
Our results yield several key insights. In the matrix completion task: First, removing the cross-ratings representation fusion (CRF) module (w/o CRF) leads to performance degradation across all datasets, underscoring its critical importance in capturing complex user interactions. Similarly, omitting the multi-channel cross-ratings module and global hypergraph learning (w/o C-R and w/o Glo) also results in reduced performance. Notably, the cross-ratings module significantly enhances results for Amazon, where the rating distribution is imbalanced, while hypergraph learning offers particular benefits for denser datasets such as ML-100K and ML-1M. The variant without contrastive learning (w/o CL) demonstrates the poorest performance, highlighting the crucial role of contrastive learning in enriching node representations and improving model robustness. Additionally, removing the balanced cross-entropy loss (w/o bCE) adversely affects performance, further confirming its effectiveness in capturing correlations between closely-rated scores. Furthermore, our analysis indicates that the cross-ratings module (C-R) has the most substantial impact on the loss function, whereas node relation regularization (NRR) exhibits the least effect, aligning with our parameter sensitivity experimental results. Finally, while the ordinal cross-entropy loss (with OCE) performs similarly to the balanced cross-entropy loss, the latter is preferred due to its slight performance edge and superior computational efficiency, making it a more practical choice for implementation.
In the recommendation task: First, removing all contrastive learning mechanisms (w/o CL) leads to a significant drop in performance, highlighting the crucial role of contrastive learning in enhancing recommendation accuracy. Second, the absence of the multi-channel cross-rating contrastive learning module (w/o C-R) also substantially degrades model performance, especially on long-tail and sparse datasets such as Amazon, indicating its effectiveness in addressing data sparsity and imbalanced distributions. Third, eliminating the global hypergraph structure learning module (w/o Glo) has the greatest impact on dense datasets, suggesting its importance in capturing higher-order dependencies and improving model generalization. Fourth, the removal of the cross-rating fusion module (w/o CRF) results in a consistent decrease in recommendation performance across all datasets, confirming the benefit of information fusion in modeling complex user behaviors. Fifth, the design of the loss function is equally vital; replacing the original balanced cross-entropy with either the standard or ordinal variant (w/o bCE, with OCE) results in noticeable performance degradation. Finally, the ablation of the node relation regularization (w/o NRR) shows minimal impact on overall performance.
5.6. Impact of imbalanced data distribution (RQ5)
@@ 5.6.1. Impact of Imbalanced User Interactions
We evaluate MHCL’s performance across users with varying interaction sparsity by dividing them into three groups: ‘Inactive’ (bottom 80% by interaction frequency), ‘Normal’ (middle 15%), and ‘Active’ (top 5%). We then compare MHCL with two leading baselines on two datasets, with results shown in Table 6. MHCL consistently outperforms the baselines, with the largest gains observed among Inactive users. This improvement is driven by InfoNCE loss, which mitigates popularity bias by enhancing representation uniformity.
@@ 5.6.2. Impact of Imbalanced Rating Interactions
Experimental results according to different rating categories are shown in Table 6. Notably, ratings 3, 4, and 5 dominate in both the Amazon and ML-100K datasets, while ratings 1 and 2 are significantly sparser (less than 10% in Amazon, 20% in ML-100K). This finding highlights significant improvements in long-tail ratings (‘1’ and ‘2’) without sacrificing performance in the head classes, which underscores the effectiveness of our MHCL in reshaping embedding distributions to better capture long-tail characteristics. While most methods perform well on the majority ratings, our approach significantly outperforms others on the much sparser ratings 1 and 2, demonstrating its effectiveness in handling data sparsity and long-tail distributions.
Table 6. Impact of imbalanced data distribution in Amazon and ML-100K datasets.
Dataset Method Each user group Each rating group
Inactive Normal Active 1-point 2-point 3-point 4-point 5-point
Amazon GIMC 0.8910 0.7477 0.6262 5.3752 3.2765 1.2233 0.2355 0.4482
RGCL 0.8856 0.7465 0.6238 5.1629 3.1658 1.2189 0.2364 0.4486
MHCL 0.8751 0.7381 0.6234 4.6893 2.9268 1.2170 0.2380 0.4480
ML-100K GIMC 0.8275 0.7615 0.7358 2.9593 1.4625 0.3947 0.2769 1.1116
RGCL 0.8266 0.7622 0.7360 2.9850 1.4698 0.3944 0.2790 1.1005
MHCL 0.8074 0.7484 0.7317 2.7040 1.3690 0.3986 0.2790 1.0894
5.7. Parameter Sensitivity (RQ6)
5.7.1. Effect of Hyperparameters $\alpha$, $\beta$ and $\lambda$.
We analyze the sensitivity of parameters $\alpha$, $\beta$, and $\lambda$ on contrastive learning (CL) losses by conducting parameter sensitivity experiments with $\alpha$, $\beta$, and $\lambda$ on Amazon and ML-100K datasets, using values from {0.001, 0.005, 0.01, 0.05, 0.1}. Results for $\alpha$, $\beta$ and $\lambda$ are shown in Figures LABEL:alpha_amazon, LABEL:alpha_ml, LABEL:beta_amazon, LABEL:beta_ml, LABEL:lambda_amazon and LABEL:lambda_ml. The detailed settings of $\alpha$, $\beta$, and $\lambda$ for each dataset are as follows: (1) Yelp: $\alpha$ = 0.05, $\beta$ = 0.01, $\lambda$ = 0.01; (2) Amazon: $\alpha$ = 0.05, $\beta$ = 0.01, $\lambda$ = 0.01; (3) ML-100K: $\alpha$ = 0.01, $\beta$ = 0.01, $\lambda$ = 0.01; (4) ML-1M: $\alpha$ = 0.01, $\beta$ = 0.01, $\lambda$ = 0.01; (5) YahooMusic: $\alpha$ = 0.01, $\beta$ = 0.01, $\lambda$ = 0.005; (6) Douban: $\alpha$ = 0.05, $\beta$ = 0.01, $\lambda$ = 0.01; (7) CD: $\alpha$ = 0.05, $\beta$ = 0.01, $\lambda$ = 0.01; (8) Alibaba: $\alpha$ = 0.01, $\beta$ = 0.01, $\lambda$ = 0.01. Increasing these parameters from 0 to 0.01 improves MHCL’s performance, indicating that contrastive learning enhances node representation learning. However, excessively large values diminish the model’s main loss function, negatively impacting performance. The optimal range for $\alpha$ and $\beta$ is between 0.01 and 0.05, highlighting the need to balance the main loss with contrastive learning for optimal results. Moreover, despite having fewer users and items, MHCL requires more hyperedges to capture complex high-order correlations.